OpenStudy (anonymous):

Solve for x: |x + 2| + 16 = 14

OpenStudy (praxer):

no solution exist. as Ix+2I is negative.

OpenStudy (phi):

the first step is add -16 to both sides, like this |x + 2| + 16 -16 = 14-16 simplify to get |x + 2| = -2 notice that the absolute value of anything is positive. in other words, the left side will be 0 or bigger. Never negative.

OpenStudy (anonymous):

oh ok, that's what was confusing me. thank you